OpenStudy (anonymous):

FAN AND MEDAL HELP PLEASE Which set of dimensions is possible for a rectangular prism with volume 420 cm3? A. 4 cm by 8 cm by 12 cm B. 4 cm by 10 cm by 12 cm C. 5 cm by 7 cm by 12 cm D. 5 cm by 8 cm by 12 cm

OpenStudy (anonymous):

@iGreen @IrishBoy123

OpenStudy (igreen):

All 3 numbers have to get 420 when you multiply.

OpenStudy (igreen):

@familiabarajas

OpenStudy (igreen):

4 * 8 * 12 = ? 4 * 10 * 12 = ? 5 * 7 * 12 = ? 5 * 8 * 12 = ?

OpenStudy (anonymous):

tankyou

OpenStudy (anonymous):

so it would be c 5*7*12